# No Train but Gain: Language Arithmetic for training-free Language Adapters enhancement

### v1

@@ Abstract
Modular deep learning is the state-of-the-art solution for lifting the curse of multilinguality, preventing the impact of negative interference and enabling cross-lingual performance in Multilingual Pre-trained Language Models. However, a trade-off of this approach is the reduction in positive transfer learning from closely related languages. In response, we introduce a novel method called language arithmetic, which enables training-free post-processing to address this limitation. Inspired by the task arithmetic framework, we apply learning via addition to the language adapters, transitioning the framework from a multi-task to a multilingual setup. The effectiveness of the proposed solution is demonstrated on three downstream tasks in a MAD-X-based set of cross-lingual schemes, acting as a post-processing procedure. Language arithmetic consistently improves the baselines with significant gains in the most challenging cases of zero-shot and low-resource applications. Our code and models are available at https://github.com/mklimasz/language-arithmetic.
@@ 1 Introduction
The recent progress of large language models has led to a huge interest in how well they can perform not just in English but across multiple languages and invoked the rise of Multilingual Pre-trained Language Models (MLLMs) [11, 43, 2]. These models serve as general-purpose solutions that can be adapted and applied to various Natural Language Processing tasks. Notably, MLLMs demonstrate zero-shot cross-lingual capabilities, allowing them to generalise effectively to downstream tasks even when trained in a language different from the target language.
The positive transfer of abilities from both related languages and high-quality training data from unrelated languages has meant that MLLMs have reported state-of-the-art performance in low-resourced languages [24]. However, this benefit does not always extend to high-resourced languages [21]. In such cases, the quality of MLLMs tends to decrease compared to their monolingual counterparts [25, 23, among others] due to negative interference phenomena [42]. Additionally, the curse of multilinguality [11] reveals the existence of a trade-off between language coverage and model capacity. Consequently, MLLMs must carefully limit the number of languages included during the pre-training phase.
Modular deep learning (MDL) [32] methods come to the rescue to avoid the abovementioned limitations, enabling the extension of MLLMs to support any number of languages [6, 40, 34, 29, 31]. MDL methods adapt the model to arbitrary tasks and languages by isolating components from each other (and the backbone MLLM) via parameter-efficient extensions. Examples of parameter-efficient modules are LoRA [17] and adapters [37, 16], which are low-budget (in terms of parameters) bottleneck layers that increase an MLLM size by just a fraction. Language adapters especially allow the modularisation of language-specific knowledge by training on a raw, unlabelled corpus in a target language.
The limitation of the MDL and language adapters is their isolation. While they lift the curse of multinguality and prevent negative interference, at the same time, language adapters limit the possible impact of positive transfer. Previous attempts to address these challenges — such as training bilingual [26] or language-family [9] adapters — do not scale effectively. In our work, we tackle this limitation as a post-processing step. Leveraging recent insights from task arithmetic [19], specifically learning via addition, we augment language adapters with missing related language knowledge – a concept we term language arithmetic. Remarkably, this training-free approach enhances language adapters across various scenarios, from zero-shot settings to low-resource regimes.
@@
  • We conduct a cross-lingual evaluation on three downstream tasks and Multilingual Pre-trained Language Models with test cases that include zero-shot and low-resource setups.
  • We provide an analysis of language arithmetic internal components.
@@ 2 Background
Our research builds upon the contributions of Ilharco et al. [19] and Zhang et al. [45]. The following Section provides the background and serves as a gentle introduction to the concept of task vectors and task arithmetic.
@@ 2.1 Task Vectors & Task Arithmetic
Task arithmetic allows us to forge a multi-task model from a separate, task-specific set of fine-tuned models, preserving high accuracy (although a shared pre-trained starting point is required, e.g. the same Language Model). Moreover, vectors from different tasks are typically close to orthogonal, and Ilharco et al. [19] speculate that this enables the combination of task vectors via addition with minimal interference.
In the Parameter-Efficient Fine-tuning (PEFT) regime, we can reduce a task vector to only newly introduced PEFT parameters [45]. Another way of looking at this is a task vector that is a zero-valued vector for all the parameters apart from PEFT weights, i.e. a task vector would not modify the base, pre-trained checkpoint.
@@ 3 Method
We propose language arithmetic that transitions the task arithmetic concept from a multi-task to a multilingual setup. In this Section, we describe the language arithmetic alongside its application as a training-free, post-processing step to a MAD-X cross-lingual framework [29].
@@ 3.1 Language Arithmetic
We formulate a language arithmetic (LA) concept by substituting the task in task vectors and arithmetic with a language. This approach means that instead of merging downstream tasks, we target a problem of cross-lingual performance. We propose to apply learning via addition to languages, and in Section 5.3, we demonstrate the discrepancies when comparing language and task vectors. Our study focuses specifically on the language adapters [34, 29]. Due to overlapping abbreviations, we use the LA exclusively as the former, i.e. language arithmetic. In the learning via addition, we limit the parameters to language adapters and simplify the notation that $\theta$ represents the adapters’ weights and $\tau$ is referred to as a language vector. As we operate in a language space instead of a task, the $t$ is replaced with a language, i.e. its language code in the notation. The example equation describes a language arithmetic operation between an English and a Spanish adapter.
@@
Throughout the paper, the equation above is abbreviated as a function: $LA(en,es)$ with lambda as a default parameter. Whenever we use language arithmetic in a zero-shot context, we add a subscript to the function name, i.e. $LA_{0s}(en,es)$ means that language arithmetic between English and Spanish is applied to a different language, e.g. French.
@@ 3.2 Application
We apply the language arithmetic as an enhancement method for language adapters. We assume two scenarios: (i) zero-shot case, where we do not have a language adapter in the desired language and (ii) “regular” enhancement case, where we improve the existing language adapter (in high- and low-resourced languages).
We evaluate our post-processing method as an extension of the MAD-X framework [29] to challenge our solution in a cross-lingual manner. The overview of the schema is presented in Figure 1.
[…]
The MAD-X consists of the following:
@@
[…]
@@
[…]
In our work, we introduce an additional step:
  1. 5.
[…]
3.2.1 MLLM
The first step usually assumes the availability of a Multilingual Pre-trained Language Model. In our work, we focus on mBERT111bert-base-multilingual-cased [12] and XLM-R222xlm-roberta-base [11].
3.2.2 Training language & task adapters
The following two steps depend on the downstream dataset. In the most extreme case, we might have a dataset in one specific language and need to take advantage of cross-lingual performance. Therefore, step number 2 trains a language adapter on raw corpora corresponding to the task language using regular masked language modelling loss in a self-supervised manner. During this step, the MLLM is frozen. The third step freezes both the backbone and language adapter and trains a task adapter on a downstream task dataset. Given a set of tasks or if a new task appears, we can repeat this step as long as a required language adapter exists, i.e. a language adapter that matches the task’s language.
3.2.3 Additional language adapters
The last, fourth step is the heart of the MAD-X framework. We can now train language adapters in any language, leverage the pre-trained task adapter(s) and obtain a cross-lingual performance by connecting a new language adapter with a task adapter (i.e. routing first via language adapter and then task) to obtain language-task pair unseen during training. The growing pool of pre-trained adapters can be accessed at public repositories like AdapterHub [28] and reused for further use cases.
@@ 3.2.4 Post-processing via Language Arithmetic
First, we assume the use-case where the pool of language adapters does not contain the target languages, i.e. zero-shot scenario. In this study, we add English as a task-specific language to a language related to the target. In the second language adapter enhancement scenario, we apply language arithmetic in its default form, where the goal is to enhance the existing target language adapter (via either English or a related language). At last, we investigate (simulated and actual) low-resource cases.
@@ 4.1 Experimental setup
We use the AdapterHub library [28] for all our experiments. For language adapters, we train on the Wikipedia corpora33320331101.xx checkpoint https://huggingface.co/datasets/wikimedia/wikipedia for 250k steps with a learning rate of 1e-4, an effective batch size of 64 using a single GPU [^4] Nvidia A100/RTX3090/V100 and the same initialisation. For task-specific training, we train for 100 epochs with the same learning rate and a batch size set to 16. We choose the final checkpoint based on validation dataset performance (for language adapters, we evaluate on a held-out subset of Wikipedia).
Downstream evaluation is performed on three tasks: Named Entity Recognition (NER), Natural Language Inference (NLI) and Question Answering (QA), covering jointly 13 languages555ar, bg, de, el, es, fr, hi, ru, sw, tr, ur, vi, zh; XQuAD does not cover 4 languages (bg, fr, sw, ur), while the training - to perform cross-lingual evaluation - is performed on the English data. For the NER task, we use the WikiANN [35] dataset and for NLI - XNLI [10]. The QA evaluation is done on XQuAD [5] (we split data 50/50 into valid/test datasets), and the training uses SQuAD 1.1 [36]. Additionally, to evaluate a low-resource scenario for a language not covered during MLLM pre-training, we leverage the Assamese subset from IndicXNLI [1].
@@ 4.2 Zero-shot evaluation
Table 1: Zero-shot results of language arithmetic (LA) per language, where one side of the arithmetic equation is the English adapter, and the other is related to the evaluated language adapter (e.g. Italian for Spanish evaluation). As baselines, we report using both adapters in isolation (pEA - proxy via English adapter, pRA - proxy via a related language adapter).
ar bg de el es fr hi ru sw tr ur vi zh AVG
NER (F1 score)
mBERT pEA 33.45 75.62 78.10 69.80 73.42 78.25 63.03 61.63 63.14 72.56 28.72 68.81 36.99 61.81
pRA 33.54 74.51 76.76 69.70 71.53 78.37 58.01 62.05 61.85 68.71 25.08 67.44 33.30 60.07
$LA_{0s}(en,REL)$ 37.22 75.83 77.89 73.42 73.42 78.51 64.11 62.95 64.80 72.56 30.39 69.25 37.65 62.77
[…]
pRA 39.07 75.24 73.12 65.00 67.01 72.74 66.43 58.14 57.53 64.60 36.72 55.82 18.91 57.72
$LA_{0s}(en,REL)$ 43.19 76.57 74.27 70.58 73.48 74.75 68.64 60.37 63.02 69.45 37.22 60.68 23.44 61.20
NLI (Accuracy)
mBERT pEA 61.66 65.43 68.40 62.08 72.22 71.84 58.34 65.87 46.25 56.93 56.49 68.46 66.73 63.13
pRA 53.87 64.49 66.43 61.14 70.84 70.68 54.47 63.21 43.33 54.43 54.51 67.43 66.63 60.88
$LA_{0s}(en,REL)$ 62.06 65.23 68.42 62.44 71.52 72.20 57.96 65.19 46.51 56.95 56.13 69.28 67.33 63.17
[…]
pRA 67.35 76.01 73.67 70.56 69.92 71.46 67.70 74.05 55.89 68.28 62.93 72.36 71.30 69.34
$LA_{0s}(en,REL)$ 68.66 76.69 73.81 73.03 76.97 76.11 68.74 73.65 60.12 70.16 63.87 72.51 73.15 71.35
QA (F1 score)
mBERT pEA 55.93 - 71.93 51.46 72.39 - 50.65 67.46 - 44.99 - 63.51 54.12 59.16
pRA 41.76 - 71.14 53.03 71.67 - 43.22 66.31 - 43.68 - 65.05 54.97 56.76
$LA_{0s}(en,REL)$ 55.93 - 71.32 53.91 72.61 - 50.40 66.36 - 44.64 - 64.16 54.46 59.31
[…]
pRA 62.54 - 69.58 64.98 72.60 - 63.00 69.09 - 63.63 - 70.07 59.15 66.07
$LA_{0s}(en,REL)$ 64.87 - 72.69 71.69 74.45 - 64.01 70.11 - 64.98 - 71.55 61.49 68.43
[…]
Table 1 presents the results of the zero-shot experiment. Language arithmetic consistently outperforms the proxy baselines, reaching up to an improvement of over $3$ F1 points in the XLM-R NER setup and almost $1.5$ F1 for QA. Additionally, we looked at the $\lambda$ parameter derived on a validation dataset. While most cases set the value to over $0.5$ (i.e. preferring the English side, given $LA_{0s}(en,REL)$), the preferred values did not showcase any consistency and pattern. We analyse that phenomenon in Section 5.1.
4.3 Language Arithmetic evaluation

### v2

@@ Abstract
Modular deep learning is the state-of-the-art solution for lifting the curse of multilinguality, preventing the impact of negative interference and enabling cross-lingual performance in Multilingual Pre-trained Language Models. However, a trade-off of this approach is the reduction in positive transfer learning from closely related languages. In response, we introduce a novel method called language arithmetic, which enables training-free post-processing to address this limitation. Extending the task arithmetic framework, we apply learning via addition to the language adapters, transitioning the framework from a multi-task to a multilingual setup. The effectiveness of the proposed solution is demonstrated on three downstream tasks in a MAD-X-based set of cross-lingual schemes, acting as a post-processing procedure. Language arithmetic consistently improves the baselines with significant gains, especially in the most challenging case of zero-shot application. Our code and models are available at https://github.com/mklimasz/language-arithmetic.
No Train but Gain:
Language Arithmetic for training-free Language Adapters enhancement
Mateusz Klimaszewski1,2, Piotr Andruszkiewicz1, Alexandra Birch2 1Warsaw University of Technology, 2University of Edinburgh Correspondence: mklimasz@ed.ac.uk
@@ 1 Introduction
The recent progress of large language models has raised the question of how well they perform not just in English but across multiple languages which has spurred interest in Multilingual Pre-trained Language Models (MLLMs) Conneau et al. (2020); Workshop (2023); Alves et al. (2024). These models serve as general-purpose solutions that can be adapted and applied to various Natural Language Processing tasks. Notably, MLLMs demonstrate zero-shot cross-lingual capabilities, allowing them to generalise effectively to downstream tasks even when pre-trained in a language different from the target language.
The positive transfer of abilities from both related languages and high-quality training data from unrelated languages has meant that MLLMs have reported state-of-the-art performance in low-resourced languages Muennighoff et al. (2023). However, this benefit does not always extend to high-resourced languages Kocmi et al. (2021). In such cases, the quality of MLLMs tends to decrease compared to their monolingual counterparts (Nozza et al., 2020; Martin et al., 2020, among others) due to negative interference phenomena Wang et al. (2020). Additionally, the curse of multilinguality Conneau et al. (2020) reveals the existence of a trade-off between language coverage and model capacity. Consequently, MLLMs must carefully limit the number of languages included during the pre-training phase.
Modular deep learning (MDL) Pfeiffer et al. (2023) methods help to avoid negative interference and limited model capacity , enabling the extension of MLLMs to support any number of languages Bapna and Firat (2019); Üstün et al. (2020); Philip et al. (2020); Pfeiffer et al. (2020b, 2022). MDL methods adapt the model to arbitrary tasks and languages by isolating components from each other (and the backbone MLLM) via parameter-efficient extensions. Examples of parameter-efficient modules are adapters Rebuffi et al. (2017); Houlsby et al. (2019), which are low-budget (in terms of parameters) bottleneck layers that increase an MLLM size by just a fraction. Language adapters Pfeiffer et al. (2020b) allow the modularisation of language-specific knowledge by training on a raw, unlabelled corpus for specific languages.
The limitation of the MDL and language adapters is their isolation. While they lift the curse of multinguality and prevent negative interference, at the same time, language adapters limit the possible impact of positive transfer. Previous attempts to address these challenges — such as training bilingual Parović et al. (2022) or language-family Chronopoulou et al. (2023b) adapters — do not scale effectively. In our work, we tackle this limitation as a post-processing step. Leveraging recent insights from task arithmetic Ilharco et al. (2023), specifically learning via addition, we augment language adapters with missing related language knowledge – a concept we term language arithmetic. Remarkably, this training-free approach can enhance not only existing language adapters but also offer zero-shot performance.
@@
  • We conduct a cross-lingual evaluation on three downstream tasks (NER, NLI and QA) and two Multilingual Pre-trained Language Models (XLM-R, mBERT) with test cases that include zero-shot and low-resource setups in a diverse group of 13 languages.
  • We provide an analysis of language arithmetic internal components (including a comparison with task arithmetic) and show improvement up to 3 F1 points without any additional training involved.
@@ 2 Background
Our research builds upon the task arithmetic contributions of Ilharco et al. (2023) and Zhang et al. (2023). The following Section provides the background and serves as a gentle introduction to the concept of task vectors and task arithmetic.
@@ 2.1 Task vectors & Task arithmetic
Task arithmetic allows us to forge a multi-task model from a separate, task-specific set of fine-tuned models, preserving high accuracy (although a shared pre-trained starting point is required, e.g. the same Language Model). Moreover, vectors from different tasks are typically close to orthogonal, and Ilharco et al. (2023) speculate that this enables the combination of task vectors via addition with minimal interference.
In our work, we focus on parameter-efficient fine-tuning (PEFT) and use language adapters. Therefore, we reduce the task vector and underlying model weights represented by $\theta$ to newly introduced parameters (i.e. we exclude the backbone MLLM, which is frozen across all the models, following the work of Zhang et al. (2023))
@@ 3 Method
We propose language arithmetic that transitions the task arithmetic concept from a multi-task to a multilingual setup. In this Section, we describe the language arithmetic alongside its application as a training-free, post-processing step to a MAD-X cross-lingual framework Pfeiffer et al. (2020b).
@@ 3.1 Language arithmetic
We formulate a language arithmetic (LA) concept by substituting the task in task vectors and arithmetic with a language. This approach means that instead of merging downstream tasks, we target a problem of cross-lingual performance. We propose to apply learning via addition to languages, and in Appendix E, we demonstrate the discrepancies when comparing language and task vectors. Our study focuses specifically on the language adapters Philip et al. (2020); Pfeiffer et al. (2020b). Due to overlapping abbreviations, we use the LA exclusively as the former, i.e. language arithmetic. In the learning via addition, we limit the parameters to language adapters and simplify the notation that $\theta$ represents the adapters’ weights and $\tau$ is referred to as a language vector. As we operate in a language space instead of a task, the $t$ is replaced with a language, i.e. its language code in the notation. The example equation describes a language arithmetic operation between an English and a Spanish adapter.
@@
Throughout the paper, the equation above is abbreviated as a function: $LA(en,es)$ with lambda as a default parameter. Additionally, for clarity reasons, we denote target language in a subscript to distinguish different use-cases. For example, $LA_{fr}(en,es)$ means that language arithmetic between English and Spanish is applied to a different language - in this case French (zero-shot application), or $LA_{es}(en,es)$ meaning that the target is Spanish (non-zero-shot use-case, due to Spanish being a part of the LA equation).
Figure 1: Language arithmetic as an extension of the MAD-X framework. Given language and task adapters (left), language arithmetic (right) enables post-processing, training-free improvement in two use-cases: (i) zero-shot where a language adapter for a target language was not trained (presented in the figure as Spanish, which was not part of existing language adapters pool, $LA_{es}(en,fr)$) or (ii) to improve existing language adapters via arithmetic with either related language or a language on which task adapter was trained (e.g. $LA_{fr}(en,fr)$).
@@ 3.2 Application
We evaluate our post-processing method as an extension of the MAD-X framework Pfeiffer et al. (2020b) to challenge our solution in a cross-lingual manner. The overview of the schema is presented in Figure 1.
The MAD-X consists of the following steps:
@@
    Cross-lingual inference
In our work, we introduce an additional step:
@@
    Post-processing via language arithmetic
In the following Sections 3.2.1-3.2.4, we present the framework and our proposed post-processing via language arithmetic extension, exploring two use cases: (i) a zero-shot case, where a target language adapter does not exist and (ii) an enhancement case, where we prove existing language adapters (in high- and low-resourced languages).
3.2.1 Training language adapter(s)
In the first step, the MAD-X framework trains language adapters. These adapters are trained on raw corpora using masked language modelling loss in a self-supervised manner. The MLLM is frozen during this step, and we only optimise the newly introduced adapter. The training must be done for languages corresponding to the downstream tasks (e.g. if we have an English NER dataset, we need an English language adapter, apart from other desired target languages). Additionally, the adapters form a pool that is leveraged during cross-lingual inference.
3.2.2 Training task adapters
The following step freezes a backbone MLLM and a language adapter and trains a task adapter on a downstream task dataset. Given a set of tasks or if a new task appears, we can repeat this step as long as the required language adapter exists in the available pool, i.e. a language adapter that matches the task’s language.
3.2.3 Cross-lingual inference
Having trained a task adapter, we can leverage a pool of pre-trained language adapters and obtain a cross-lingual performance by connecting any existing language adapter with a newly trained task adapter (i.e. routing first via language adapter and then task adapter). The growing pool of pre-trained adapters can be accessed at public repositories like AdapterHub Pfeiffer et al. (2020a) and reused for further use cases.
@@ 3.2.4 Post-processing via language arithmetic
Our method builds upon the MAD-X framework in two enhancement scenarios.
First, we assume a situation where the pool of language adapters does not contain a desired target language, i.e. a zero-shot scenario. In contrast to the previous works that try to improve existing adapters, this use-case is an alternative to routing via either a related language or a task language (here, by task language, we understand the language on which the task adapter was trained, in contrast to a target language - on which we want to evaluate). Instead of choosing the better-performing proxy, language arithmetic proposes to combine these two (with better results, as shown in Section 4.2).
In the second language adapter enhancement scenario, we apply language arithmetic as a more common goal, trying to improve existing language adapters; however, our method does that without any training. Here, we combine the existing target language adapter with either a related language (we define related languages in Section 4.1.1) or, once again, a task language.
@@ 4.1 Experimental setup
4.1.1 Datasets
Downstream evaluation is performed on three tasks: Named Entity Recognition (NER), Natural Language Inference (NLI) and Question Answering (QA), covering jointly 13 languages111ar, bg, de, el, es, fr, hi, ru, sw, tr, ur, vi, zh; XQuAD does not cover 4 languages (bg, fr, sw, ur), while the training - to perform cross-lingual evaluation - is performed on the English data. For the NER task, we use the WikiANN Rahimi et al. (2019) dataset and for NLI - XNLI Conneau et al. (2018). The QA evaluation is done on XQuAD Artetxe et al. (2020) (we split data 50/50 into valid/test datasets), and the training uses SQuAD 1.1 Rajpurkar et al. (2016). Additionally, to evaluate a low-resource scenario for a language not covered during MLLM pre-training, we leverage the Assamese subset from IndicXNLI Aggarwal et al. (2022).
Related languages
To automatically establish a related language needed for language arithmetic, we query URIEL and lang2vec library Littell et al. (2017); Malaviya et al. (2017). During the related language query, we limited the options to 13 downstream task languages for which we had already pre-computed language adapters. This limitation means that for some languages, we would be able to find a stronger performing pairing and that the ceiling for our method is higher than the presented (we denote that the performance of hypothetical paring would also depend on data availability, i.e. a paired language must be not only related but also have representative corpora; based on this we show analysis and improved performance in Section 5.2). However, considering the limitations of the lang2vec, we decided to keep this simplification. At last, a language can have a set of equally good related languages. Therefore, in practical terms, it is not feasible for our study to train all possible options for each language - our simplification stands as a reasonable, real-world proxy. We provide the list of related languages in the Appendix A.
4.1.2 Implementation & training
In our work, we focus on two of the most popular multilingual PLMs [^2] According to downloads from Huggingface hf.co/models?language=multilingual&sort=downloads: mBERT333bert-base-multilingual-cased Devlin et al. (2019) and XLM-R444xlm-roberta-base Conneau et al. (2020). We implement our method using the AdapterHub library Pfeiffer et al. (2020a). For language adapters, we train on the Wikipedia corpora55520331101.xx checkpoint hf.co/datasets/wikimedia/wikipedia for 250k steps with a learning rate of 1e-4, an effective batch size of 64 using a single GPUand the same initialisation. For task-specific training, we train for 100 epochs with the same learning rate and a batch size set to 16. We choose the final checkpoint based on validation dataset performance (for language adapters, we evaluate on a held-out subset of Wikipedia). In our main experiments, we report the scores as an average over three independent runs with different seeds (for both language and task adapters). Additionally, to improve efficiency and reduce GPU memory utilisation, we adopt a half-precision (FP16) setting.
@@ 4.2 Zero-shot evaluation
The zero-shot evaluation assumes a scenario where the language adapter pool does not contain a desired target language (e.g. lack of Spanish in Figure 1). The baselines are based on routing, i.e. we proxy either by an English adapter (proxy via a task language, as the task adapter was trained using English data) or a related language (e.g. French for Spanish). Language arithmetic serves a solution that, instead of choosing a better proxy, combines the adapter’s tuple: $LA_{t}(en,rel)$, where $rel$ symbolises a related language and $t$ stands for a target language (e.g. $LA_{es}(en,fr)$).
Figure 2: Zero-shot XLM-R language arithmetic evaluation, where one side of the arithmetic is an English adapter, and the other is related to the target language adapter (e.g. French for Spanish - $LA_{es}(en,fr)$). The values above bars present a relative difference to a better proxy. See Figure 6 for the mBERT model.
Figures 2 and 6 present the results of the zero-shot experiment. Language arithmetic consistently outperforms the proxy baselines for all the setups, reaching over $3.1$ F1 points improvement in the NER task and $1.1$ F1 for QA (XLM-R). These results indicate that language arithmetic is a feasible, low-cost method that one can apply in the lack of an existing target language adapter.
Additionally, we investigate how the $\lambda$ parameter impacts the downstream evaluation. The goal was to understand how much weight is given to English vs related language. We looked at the validation performance over different $\lambda$ thresholds. While most cases set the value to over $0.5$ (i.e. preferring the English side, given $LA_{t}(en,rel)$), the preferred values did not showcase any consistency and pattern. We analyse this further in Section 5.1.
4.3 Improving existing language adapters
This evaluation assumes that a target language adapter exists in the adapter pool. We test two cases, i.e. $LA_{t}(en,t)$ and $LA_{t}(rel,t)$, where $rel$ is once again a related language and $t$ is the target language. Additionally, we provide a combination of these two approaches (referred to as $LA_{t}(en/rel,t)$), where for each language, we choose a better pairing (so either $en$ or $rel$). This solution resembles a practical compromise between cost and performance and serves as a proxy for the ceiling of our method (discussed in Section 4.1.1).
The results are presented in Figures 3 and 7. Compared to the baseline direct application of a target language adapter (i.e., the MAD-X method), the gains are not as significant as in the case of the zero-shot scenario. Moreover, in contrast to the previous Section’s study, MLLMs showcase a different behaviour, as language arithmetic provides less benefit for XLM-R than for the mBERT model (e.g. $+0.38$ XLM-R vs $+2.41$ mBERT in F1, QA).